OpenStudy (anonymous):

Help ?? Harry has a jar that contains 75% of apple juice and the rest water. The bottle has 1 liter of water. Part A: Write an equation in one variable that can be used to find the total number of liters of apple juice and water in the jar. Define the variable used in the equation. (5 points) Part B: How many liters of apple juice are present in the jar? Show your work. (5 points)

OpenStudy (anonymous):

Well you know apple juice = 75% so water = 25%

OpenStudy (anonymous):

theres 3 liters of juice

OpenStudy (anonymous):

I need help setting up the equation :/

OpenStudy (anonymous):

Right.

OpenStudy (anonymous):

Let's say W = water and J = juice.

OpenStudy (anonymous):

x = liters

OpenStudy (anonymous):

Since there's 1 liter of water, we can write xW, and 3 liters of juice, we can write 3xJ. So our equation is xW + 3xJ

OpenStudy (anonymous):

Define the variable used in the equation. Well you know apple juice = 75% so water = 25% Keep in mind that percentage is the number over 100. So when you have 75% of apple juice, it means you have 75/100 of apple juice from the total mixture of apple juice and water. The same applies for 25% water. Part B: How many liters of apple juice are present in the jar? Have x(or anything you want) represent the total volume of mixture. So you have for apple juice: 75/100 * x water: 25/100 * x You know the water is 1 liter. So you have 25/100 * x = 1 This is how you set up the equations. Can you do it from here?

OpenStudy (anonymous):

@Reaper534 Google Much?

OpenStudy (anonymous):

it was the simples example

OpenStudy (anonymous):

Alright i can only have 1 variable tho ... so would i do 75x+25 ?

OpenStudy (anonymous):

You can just do x + 3x.

OpenStudy (anonymous):

Be sure to say 'x' represents water and '3x' represents Juice.

OpenStudy (anonymous):

total= xW+3xJ like green said that means that the total is equal to 1 part water and three parts juice. You are given water = 1, so plugging in one for x, y=x+3x y=1+3 y=4

OpenStudy (anonymous):

@Reaper534 She already knows how much Juice there is in the jar(3 liters), all she has to do is add that to the 1 liter of water. 3 + 1 = 4.

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

I got it from here , thanks :)